# 3ICREA, Catalonia, Spain

### v1

@@ Introduction
This paper describes a multifaceted analysis of training datasets pertinent to chronic diseases aimed at uncovering potential discrepancies that could lead to biases in the resulting ML models. Our research substantiates the premise that demographic parity within datasets does not inherently ensure uniformity in algorithmic performance. That is to say, even datasets that are ostensibly equitable in terms of demographic attributes may still yield models with performance discrepancies. Initiating our analysis with a common ML performance metric, the Area Under the Receiving Operating Characteristic curve (AUROC or AUC), we measure the predictive efficacy of the models. Subsequently, our examination extends to more profound dataset attributes impacting model behavior, particularly data complexity and systematic arbitrariness. Our methodology provides a comprehensive approach for the assessment of training data from the perspective of algorithmic fairness. To the best of our knowledge, this investigation is the first to test systematic model arbitrariness in the healthcare domain.
@@ Discussion
Within the healthcare domain, the legal and ethical dimensions of decision-making are of paramount importance 48. The findings of this study highlight some characteristics of model performance that are not typically reported, but that hold considerable potential to influence clinical practice. Specifically, systematic arbitrariness in model outputs could undermine clinician confidence in ML and diminish the acceptability of such models. We propose datasets are tested for systematic arbitrariness before being used in clinical settings. In nearly half of the datasets we studied, older patients with chronic diseases face the risk of health inequities 49; 50 due to data that is suboptimal for modeling their health outcomes as compared to younger patients.
@@ Acknowledgments
https://anonymous.4open.science/r/health_disparities_Chronic_Diseases-32D4/README.md

### v2

@@ Introduction
This paper describes a multifaceted analysis of training datasets pertinent to chronic diseases aimed at uncovering potential discrepancies that could lead to biases in the resulting ML models. Our research substantiates the premise that demographic parity within datasets does not inherently ensure uniformity in algorithmic performance. That is to say, even datasets that are ostensibly equitable in terms of demographic attributes may still yield models with performance discrepancies. Initiating our analysis with a common ML performance metric, the Area Under the Receiving Operating Characteristic curve (AUROC or AUC), we measure the predictive efficacy of the models. Subsequently, our examination extends to more profound dataset attributes impacting model behavior, particularly data complexity and systematic arbitrariness. Our methodology provides a comprehensive and systematic framework for assessing training data from the perspective of algorithmic fairness, incorporating three key steps: (1) Model Performance Analysis, which evaluates disparities in predictive accuracy between demographic groups using robust cross-validation and bootstrapping techniques; (2) Data Complexity Metrics, leveraging established metrics to explore the intrinsic challenges of the datasets, such as feature informativeness and class imbalance; and (3) Systematic Arbitrariness Examination, quantifying inconsistencies in model predictions and their concentration across specific data subsets. To the best of our knowledge, this investigation is the first to test systematic model arbitrariness in the healthcare domain.
@@ Discussion
Within the healthcare domain, the legal and ethical dimensions of decision-making are of paramount importance 48. The findings of this study highlight some characteristics of model performance that are not typically reported, but that hold considerable potential to influence clinical practice. Specifically, systematic arbitrariness in model outputs could undermine clinician confidence in ML and diminish the acceptability of such models. In nearly half of the datasets we studied, older patients with chronic diseases face the risk of health inequities 49; 50 due to data that is suboptimal for modeling their health outcomes as compared to younger patients. To address these challenges, datasets and models should undergo rigorous testing for systematic arbitrariness prior to deployment, incorporating self-consistency metrics to detect and mitigate prediction discrepancies across demographic groups. Model developers must prioritize detailed fairness audits, including systematic arbitrariness and demographic-specific performance metrics, as part of standard validation protocols. Healthcare practitioners are integral to the ethical implementation of ML models by insisting on transparency from developers, carefully evaluating outputs for potential biases, and complementing ML recommendations with clinical expertise. They should actively promote inclusivity by pinpointing data deficiencies that disadvantage certain groups and collaborating with developers to mitigate these gaps, while encouraging continuous model evaluation to refine performance over time.
@@ Acknowledgments
Ethics and Impact Statement
This study did not involve new data collection; all accessed and processed data were anonymized in compliance with ethical standards and protecting participant privacy.
@@
https://github.com/IoannisBil/health_disparities_Chronic_Diseases/blob/main/README.md